OpenStudy (anonymous):

What is an equation for the inverse variation? y varies inversely with x, and y = 2 when x = 6.

OpenStudy (anonymous):

\[y \alpha \frac{ 1 }{ x }\] \[y = \frac{ k }{ x }\] \[2 = \frac{ k }{ 6 }\] \[12 = k\] Therefore... \[y = \frac{ 12 }{ x }\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

thats ok :)